OpenStudy (anonymous):

Using a directrix of y = 2 and a focus of (3, -4), what quadratic function is created?

OpenStudy (anonymous):

f(x) = one twelfth (x - 3)2 - 1 f(x) = -one sixth (x + 3)2 + 1 f(x) = one sixth (x - 3)2 + 1 f(x) = -one twelfth (x - 3)2 - 1 HELP PLEASE??? i woudlnt post unless i really dk the material

OpenStudy (anonymous):

@wio